If a square and a rhombus stands on same base, then find the ratio between area of square and rhombus ? (For class 9)
1
Expert's answer
2019-09-27T12:54:10-0400

Area of square: "S_1=a^2"

Area of rhombus: "S_2=a^2\sin\alpha", where "\alpha" is a smaller angle of the rhombus.


"\frac{S_1}{S_2}=\frac{a^2}{a^2\sin\alpha}=\frac{1}{\sin\alpha}"



Answer:

"\frac{1}{\sin\alpha}"